Show that for small thermal expansions of a solid object the linear and 

volume coefficients of expansion are related by β ≈ 3α.



1
Expert's answer
2021-03-23T11:43:56-0400


From the linear thermal expansion law we know that

"ΔL = αL_iΔT \\

dL = αL_idT"

Note that, every dimension of the cube will expand linearly and the final volume of the cube will equal to

"V_f = L_fw_fh_f"

The initial volume of the cube before raising its temperature is

"V_i = L_iw_in_i"

From the law of volume thermal expansion we know, that:

"ΔV = βV_iΔT \\

dV = βV_idT"

Note that, the cube’s dimensions are equal; therefore,

"L_i = w_i = h_i = L"

Because the cubic is made of one material so that, the three dimensions will expand by the same rate; therefore,

"L_f=w_f=h_f=L_f"

From all the above,

"V_i=L^3_i \\

V_f = L^3_f"

Note that,

"dL = L_f -L_i \\

dL +L_i = L_f \\

V_f = (L_i+dL)^3 \\

dL = αL_idT \\

V_f = (L_i+αL_idT)^3 = (L_i(1+α))^3 \\

V_i = L^3_i \\

V_f = V_i(1+α)^3 \\

V_f = V_i (1+αdT + 2αdT+2α^2dt^2)(1+αdT) \\

V_f = V_i(1+αdT + 2αdT+2α^2dT^2 + α^2dT + α^3dT^3) \\

V_f = V_i(1+3αdT + 3α^2dT^2+α^3dT^3)"

We are supposing that the change in the length dL is very infinitesimal. This means that, "dL^2" is very close to zero and "dL^3" is also very close to zero. So that, we will neglect themselves

"V_f = V_i(1+3αdT + 0+0) \\

V_f = V_i(1+3αdT) \\

V_f = V_i+3αV_idT \\

V_f- V_i = 3αV_idT \\

dV = 3αV_idT \\

dV = βV_idT \\

βV_idT = 3αV_idT \\

β = 3α"